OpenStudy (anonymous):

a garden is shaped like a rectangle whose perimeter is 308 ft. the length is 10 times as long as the width. find the length and width

OpenStudy (anonymous):

P=l + l + w + w l=10*w P=10w+10w+w+w 308=22w w=308/22

OpenStudy (wolf1728):

perimeter = 2*L + 2*w = 308 L =10 * w Multiplying by -2 -2L = -20*w -2L +20*w = 0 Adding above equation to 2*L + 2*w = 308 22w =308 width = 14 Length = 140